OpenStudy (anonymous):

(-2, -1) is a solution of the equation 4x + 5y = -13. True or false

OpenStudy (lgbasallote):

is 4(-2) + 5(-1) = -13?

OpenStudy (amorfide):

(-2, -1) is a coordinate the standard co ordinate is this (x,y) -2=x -1=y substitute the values into your equation

OpenStudy (amorfide):

4(-2) +5(-1)=-13 is correct

OpenStudy (anonymous):

@amorfide Thanks that was my last question finally thanks..